Circular No. 7947

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002eh IN NGC 917 Independent reports have been received of the discoveries of an ap-

parent supernova on unfiltered KAIT CCD images obtained by LOTOSS (cf. IAUC 7906), via B. Swift and W. D. Li, University of California at Berkeley, on Aug. 1.5 and Aug. 3.4 UT (mag ∼ 16.4) and by T. Boles, Coddenham, England, in the course of the U.K. Nova/Supernova patrol on Aug. 2.10 (mag 16.5) and 3.00 UT with a 0.35-m reflector. SN 2002eh is located at α = 2^{h}26^{m}08^{s}.45, δ = +31^{o}54′45′′.2 (equinox 2000.0), which is 9′′.3 east and 1′′.0 north of the nucleus of NGC 917 (measurements by Li). A KAIT image taken on Feb. 1.2 UT showed nothing at this position (limiting mag ∼ 19.0). Boles adds that SN 2002eh is not present on earlier images taken by himself or on second Palomar Sky Survey blue and red images.

SUPERNOVA 2002ed IN NGC 5468 M. Salvo and P. Price, Research School of Astronomy and Astro-

physics, Australian National University (ANU), report that a spectrum (range 340–920 nm, resolution ∼ 0.4 nm) of SN 2002ed taken at the ANU 2.3-m (+ Double Beam Spectrograph) at Siding Spring Observatory on July 30.41 UT shows it to be similar to the type-IIP SN 1992H at early phases, according to the spectra published in Clocchiatti et al. (1996, A.J. 111, 1286) and shows a blue continuum with a broad Hα emission and Hβ with a P-Cyg profile. The expansion velocity measured from the blueshifted minimum of Hβ is ∼11 500 km/s, adopting the NED recession velocity of 2845 km/s for the host galaxy.

V2274 CYGNI R. J. Rudy, D. K. Lynch, S. Mazuk and C. C. Venturini, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Sciences, University of California at San Diego; and R. B. Perry, NASA Langley Research Center, report 0.8–2.5 μm spectroscopy of the nova V2274 Cyg using the Aerospace Corporation Near-Infrared Imaging Spectrograph and the Lick Observatory 3-m telescope on July 18.4 UT (∼ 370 days after discovery): “He I at 1.0830 μm is the dominant line in the spectrum. H I and He II lines are also present at moderate strength. [S VIII] at 0.9911 μm is the only coronal line detected. The FWHM of the emission lines is ∼1400 km/s. The spectrum is very red, and the H I lines indicate a reddening value of E(B–V) = 1.5, but there is no obvious emission from hot dust.”

2002 August 3 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother